OpenStudy (anonymous):

Solve. 4y = 144 A. 16 B. 36 C. –36 D. –16

OpenStudy (unklerhaukus):

4y = 144 to solve the equation for y, divide both sides of the equation by 4

OpenStudy (anonymous):

here it is i have 5 questions

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

i dont know

OpenStudy (unklerhaukus):

4y = 144 4y/4 = 144/4 =

OpenStudy (anonymous):

16

OpenStudy (anonymous):

an i right @UnkleRhaukus

OpenStudy (unklerhaukus):

how did you get that result?

OpenStudy (anonymous):

i just guess because i dont know it

OpenStudy (anonymous):

so im wrong

OpenStudy (unklerhaukus):

do you know how to simplify 4y/4 ?

OpenStudy (anonymous):

no

OpenStudy (unklerhaukus):

pretend to a moment that y stands for yaks 4y = 4 yaks 4y/4 = 4 yaks / 4 how many yaks are in each group if 4 yaks are divided into four groups ?

OpenStudy (anonymous):

omg i dont know how to do this i have 5 questions and 10 minutes to do it plzzzzz hurry

OpenStudy (anonymous):

36?